OpenStudy (anonymous):

help write an equation of a line in slope-intercept form with a slope of 3/4 and contains this point (-2, 7/2) @blurbendy

OpenStudy (blurbendy):

use the formula y - y1 = m(x - x1) where m is the slope and y1 and x1 are the points given. then get it in y = mx + b form